# ImLPR: Image-based LiDAR Place Recognition using Vision Foundation Models

### v1

@@ Abstract
LPR (LPR) is a key component in robotic localization, enabling robots to align current scans with prior maps of their environment. While VPR (VPR) has embraced VFM (VFM) to enhance descriptor robustness, LPR has relied on task-specific models with limited use of pre-trained foundation-level knowledge. This is due to the lack of 3D foundation models and the challenges of using VFM (VFM) with LiDAR point clouds. To tackle this, we introduce ImLPR, a novel pipeline that employs a pre-trained DINOv2 VFM to generate rich descriptors for LPR. To our knowledge, ImLPR is the first method to leverage a VFM to support LPR. ImLPR converts raw point clouds into RIV (RIV) to leverage VFM in the LiDAR domain. It employs MultiConv adapters and Patch-InfoNCE loss for effective feature learning. We validate ImLPR using public datasets where it outperforms SOTA (SOTA) methods in intra-session and inter-session LPR with top Recall@1 and F1 scores across various LiDARs. We also demonstrate that RIV outperforms BEV (BEV) as a representation choice for adapting LiDAR for VFM. We release ImLPR as open source for the robotics community.
@@ 1 Introduction
Place recognition, encompassing both VPR and LPR, provides an initial localization estimate when seeking to determine if a location has been previously visited, using a prior map of the location. Early learning-based methods for both modalities relied on specialized architectures trained on domain-specific datasets [1, 2, 3, 4, 5, 6]. While VPR has advanced by using VFMs [7, 8, 9, 10] pre-trained on large-scale datasets [11, 12, 13], LPR has been limited by the lack of suitable 3D foundation models and the challenges in adapting VFM to LiDAR data due to the modality difference.
Efforts applying foundation models on LiDAR data can be characterized in two ways: developing full 3D foundation models or converting LiDAR data into a format that can be used with VFM. Existing 3D foundation models [14, 15], designed for object detection or indoor scene analysis, are unsuitable for LPR due to their focus on specific tasks and the heavy computation required when working with 3D data with large sets of parameters. On the other hand, converting 3D point clouds into 2D images is a more straightforward approach for leveraging VFM for LPR. Two types of image projection are typically used to create a 2D representation from a 3D point cloud, BEV and RIV, as shown in Fig. 1. However, there exists a domain gap between the natural image (used to pre-train the VFM) and this projected LiDAR point cloud image. Adapting VFM for LPR requires this domain gap to be tackled. In this paper, our research addresses this issue.
@@
We present ImLPR, a novel pipeline for LPR that adapts DINOv2 [12], SOTA VFMs that have robust feature extraction capabilities due to their transferable representations. We utilize the RIV projection to represent the three LiDAR point cloud channels (reflectivity, range, and normal ratio) as an image. In doing so, we rely on modern LiDAR sensors having high resolution such that the RIV projection is similar to the natural images used to train DINOv2. In our work, we demonstrate that the RIV representation is more suitable for this than the BEV representation. By inserting and training lightweight adapters, we can keep most of the DINOv2 frozen while adapting it for LPR using RIV input images. Additionally, we present the Patch-InfoNCE loss which is used to drive patch-level contrastive learning. This loss enhances discriminability and robustness by ensuring consistent feature representations across corresponding RIV patches. Combining LiDAR’s geometric precision with DINOv2’s expressivity, ImLPR surpasses SOTA LPR methods. Our contributions include:
  • ImLPR is the first LPR pipeline using a VFM. By utilizing the RIV representation and carefully designed lightweight convolutional adapters, we close the domain gap between 3D LiDAR and 2D vision models. Freezing most DINOv2 layers preserves pre-trained knowledge during training, ensuring strong generalization and outperforming task‑specific LPR networks. We also demonstrate effective fine-tuning of the DINOv2 model for LPR.
@@ 2.2 Vision and 3D Foundation Models for 3D Data
VFMs [12, 13, 11] excel in VPR due to their robust feature representations [7, 9]. In particular, DINOv2 was pre-trained on massive datasets containing 140 million images. In contrast, 3D foundation models like PTv3 [14] and Sonata [15] are trained on smaller datasets with 23-139 thousand point clouds, and focus on object-level tasks, limiting their generality for outdoor LPR. Recently, researchers have attempted to apply VFM to point clouds [24, 25]. However, they focus on object detection or registration, and still rely on visual images, diverging from LPR. Instead, to adapt LiDAR data for use with image-based networks, traditional methods project point clouds into 2D formats like BEV [21] or RIV [26], and then use these images with ImageNet-pretrained CNNs such as ResNet [27]. However, compared to VFM, these methods lack the expressiveness needed for diverse outdoor scenes and are limited by their small training datasets and simple architectures. Additionally, domain differences between LiDAR images and visual images hinder the direct application of VFMs. Fine-tuning the entire model often leads to catastrophic forgetting, reducing generalization by shifting pre-trained knowledge. To address these challenges, we utilize DINOv2 with a MultiConv adapter to bridge domain differences while preserving almost the entire pre-trained knowledge within the VFM network. We adopt the RIV projection, which captures denser geometric detail and is more similar to normal visual images compared to the sparser BEV images [21, 6].
@@ 3 Methodology
3.1 Problem Definition
Traditional LPR encodes a LiDAR scan into a global descriptor, whereas ImLPR first renders the scan as an RGB-style image $\mathbf{I}\!\in\!\mathbb{R}^{H\times W\times 3}$ and then computes global representation $\mathbf{g}\!\in\!\mathbb{R}^{t}$ through the mapping function $\Omega=h\!\circ\!f$. The encoder $f:\mathbb{R}^{H\times W\times 3}\!\rightarrow\!\mathbb{R}^{H^{\prime}W^{\prime}\times C}$ ($H^{\prime}=\lfloor H/14\rfloor,\;W^{\prime}=\lfloor W/14\rfloor$) extracts patch features, which the aggregator $h:\mathbb{R}^{H^{\prime}W^{\prime}\times C}\!\rightarrow\!\mathbb{R}^{t}$ pools into descriptor $\mathbf{g}$. Metric learning optimizes $\Omega$ to ensure that if the spatial distance $\mathcal{D}(\mathbf{x}_{q},\mathbf{x}_{i})<\mathcal{D}(\mathbf{x}_{q},\mathbf{x}_{j})$, then the descriptor distance $d_{g}(q,i)<d_{g}(q,j)$. This aligns geometric proximity with embedding similarity.
@@ 3.2 Input Data Processing
where $W$ and $H$ denote the image width and height, $r_{i}$ is the range from the origin to $p_{i}$, $f_{\text{up}}$ is the upper bound of the vertical FOV (FOV), and $f$ denotes the total vertical FOV. To further enrich the image with geometric information, we include the three LiDAR channels—reflectivity, range, and normal ratio—in each pixel. The normal ratio is derived from the singular value ratio of the covariance matrix for the $k$-nearest neighboring points of $p_{i}$. Singular Value Decomposition (SVD) is applied to this covariance matrix, and the logarithmic ratio of the largest to smallest singular values encodes the normal information. This scalar effectively summarizes local geometric variations into a single channel, offering a computationally efficient alternative to multi-channel normal vectors [26]. This multi-channel image is used as input for ImLPR, as depicted in Fig. 2.
@@ 3.5 Patch-level Contrastive Learning
To encourage learning both locally (patch feature-level) as well as globally (descriptor-level) discriminative embeddings, in addition to the global embedding loss, we also train using a patch-level constrastive loss. This helps to promote more effective updates throughout the network, from low-level to high-level layers. To enable this patch-level contrastive learning, we select positive and negative patch pairs from two RIV images using their range channel information.
Positive Patch Mining: We transform two LiDAR range images into 3D point clouds using RIV coordinates and range channel data. Using known ground truth poses, with additional Iterative Closest Point (ICP) pose refinement, the point clouds are transformed into a common world frame. The transformed points from the second scan are reprojected onto the first scan’s RIV to identify matching patch pairs, as illustrated in Fig. 3. Positive pairs are selected based on pixel overlap and range channel similarity. Patches $p_{1}$ and $p_{2}$ are considered candidate positive pair if the fraction of pixels in the common valid set $\mathcal{P}_{\text{ov}}$, which have range measurements, exceeds $\rho_{\text{valid}}=0.5$ of the total patch size. For pixels in $\mathcal{P}_{\text{ov}}$, we compute the per-pixel range differences $d_{i}=r_{1,i}-r_{2,i}$. An adaptive threshold $\tau_{r}$ is computed using the median absolute deviation (MAD): $\tau_{r}=\text{median}(d_{i})+k\cdot\text{MAD}$, ensuring robustness to LiDAR measurement noise. A pair is confirmed positive if the mean range difference $\Delta_{r}=\frac{1}{|\mathcal{P}_{\text{ov}}|}\sum_{i\in\mathcal{P}_{\text{ov}}}|r_{1,i}-r_{2,i}|$ is less than $\tau_{r}$. This ensures that spatially consistent, range-coherent positive patch pairs are selected.
@@
Negative Patch Mining: When selecting negative patch pairs, we ensure that there is a large vertical and horizontal distance ($v_{dist}$ and $h_{dist}$) between the positive patches to prevent overlap. The cylindrical geometry is explicitly accounted for, ensuring that patches near the $0^{\circ}$ and $360^{\circ}$ boundaries are treated as spatially distant from the positive patches.
To supervise the patch-level constrastive learning, we propose the Patch-InfoNCE loss, a patch-level contrastive loss extending the InfoNCE loss [13]. Unlike global descriptor-based losses that operate on the global descriptor, the Patch-InfoNCE loss leverages each local patch embedding from the features $\mathbf{F_{1}},\mathbf{F_{2}}\in\mathbb{R}^{H^{\prime}W^{\prime}\times C}$ of the two adjacent RIV images. Supporting multiple positive pairs $(p_{1},p_{2})$ per image, we compute their cosine similarities. Specifically, for each positive feature pair $F_{1}^{p_{1}}$ and $F_{2}^{p_{2}}$ , we select negatives from both images: $n_{1}$ representing the hardest negatives sampled from $\mathbf{F_{1}}$ relative to $p_{2}$, and $n_{2}$ representing the hardest negatives sampled from $\mathbf{F_{2}}$ relative to $p_{1}$. The Patch-InfoNCE loss is defined as:
@@
where $\mathcal{P}$ is the positive pair set, $\tau_{l}$ is the temperature, and $s_{\text{p}}$, $s_{\text{n},j}$ are the positive and negative similarities. This loss promotes consistent local features for corresponding spatial locations in RIV images, enabling similar descriptors for LPR.
3.6 Truncated SmoothAP Loss with Global Image Descriptor
To train ImLPR with the global descriptor, we adopt the Truncated SmoothAP (TSAP) loss [4]. Unlike the triplet and contastive loss, TSAP optimizes Average Precision (AP) through a continuous approximation, focusing ranking evaluation on top candidates to reduce computation. For a query descriptor $g_{q}$ with positives in $\mathcal{P}^{+}$ and all batch descriptors in $\mathcal{D}$, TSAP is formulated as:
The final loss integrates both objectives as $\mathcal{L}_{\text{final}}=\mathcal{L}_{{P}}+\lambda\mathcal{L}_{\text{TSAP}}.$ where $\lambda$ is a weighting factor balancing local and global contributions. This dual-loss approach effectively captures fine-grained local context and optimizes global ranking, enhancing the overall LPR accuracy.
@@ 4.1 Experimental Setup
Method R01 R02 R03 T01 T02 T03 Average
@@
(R: Roundabout-O, T: Town-O), Bold: Best and Underline: Second-Best
Figure 4: Trajectory from HeLiPR-O, with red lines marking false positives (fewer red lines indicate better performance). It highlights ImLPR having fewer errors compared to SOTA methods.
@@ 4.2 Intra-session Place Recognition
Scans within 60 seconds of the query are excluded to prevent intra-session matching, and processing begins only after 90 seconds to ensure a sufficient database. The positive condition from the training strategy was applied. As shown in Table. 1 and Fig. 4, ImLPR surpasses all baselines on Roundabout-O and Town-O. LoGG3D-Net exhibits lower recall due to the domain shift between training and testing. BEVPlace++ ranks second, capturing scan contours but underperforms ImLPR in narrow alleys and small-scale environments, as ImLPR leverages RIV representations for fine-grained feature extraction. 3D sparse convolution methods like MinkLoc3Dv2 and CASSPR rank third, limited by the absence of large-scale pre-trained models. ImLPR achieves the highest R@1 and F1 score, demonstrating robust and discriminative intra-session LPR performance.
@@ 4.3 Inter-session Place Recognition
Table. 2 and Table. 3 show inter-session results (Database-Query) for Roundabout-O and Town-O, using scans from distinct sessions with temporal differences. ImLPR outperforms other methods across both environments, adapting to session-induced variations. BEVPlace++ achieves the second-highest Recall@1, but its lower F1 score indicates reduced precision. We believe this is due to the simple network consisting of 2D convolution without additional mechanisms such as attention. Furthermore, its BEV representation, projecting all data into a single plane, limits its ability to distinguish subtle session-specific variations. As shown in Fig. 5, BEVPlace++ has a lower F1-Recall curve than MinkLoc3Dv2 and CASSPR, despite higher recall. MinkLoc3Dv2 and CASSPR follow ImLPR in F1 score with inconsistent results across sequences. This variability highlights the benefit of using large-scale pretrained models for consistent results. LoGG3D-Net ranks lowest due to domain mismatches. ImLPR’s RIV representations precisely capture fine-grained details as shown in Fig. LABEL:fig:dino_riv, and its network design ensures top inter-session LPR performance.
@@
Method R01-R02 R01-R03 R02-R01 R02-R03 R03-R01 R03-R02 Average
@@
Method T01-T02 T01-T03 T02-T01 T02-T03 T03-T01 T03-T02 Average
@@
Table 4: Results in MulRan dataset
MulRan DCC (OS1-64)
Method 01 02 03
R@1 F1 R@1 F1 R@1 F1
LoGG3D-Net 0.117 0.310 0.075 0.144 0.090 0.187
MinkLoc3dv2 0.849 0.939 0.651 0.829 0.666 0.841
CASSPR 0.813 0.943 0.607 0.793 0.629 0.818
BEVPlace++ 0.811 0.937 0.585 0.766 0.639 0.813
ImLPR 0.909 0.965 0.861 0.945 0.825 0.918
4.4 Model Generalization Assessment
To assess the generality of ImLPR, we conducted intra-session LPR on the MulRan, NCLT, and HeLiPR-V without additional training. As these datasets lack a reflectivity channel, we used normalized intensity values. All experiments except for the MulRan are included in the supplementary.
We employed the DCC01-03 sequences from the MulRan for this evaluation. Although the training and test sets spatially overlap, they differ due to a four-year temporal gap, hardware variations (e.g., vehicle and OS1-64 LiDAR), and occlusions. For testing, points are projected into $1022\times 64$ RIV images and resized to $1022\times 126$ to address LiDAR vertical ray differences. As demonstrated in Table. 4 and Fig. 6, most methods show performance degradation due to these domain shifts. Nevertheless, ImLPR outperforms other methods across most metrics. MinkLoc3Dv2 ranks second in the DCC sequences by using 3D sparse convolution, which handles occlusions by avoiding feature generation in empty spaces. BEVPlace++ ranks third or fourth, as its BEV images include features for occluded or empty regions, reducing accuracy. In contrast, ImLPR mitigates the impact of low resolution and occlusion in RIV images using its powerful feature extractor. This outcome demonstrates ImLPR’s ability to generalize on the MulRan dataset.
4.5 BEV and RIV Representations with Patch-InfoNCE Loss
To evaluate the effectiveness of BEV and RIV representations for ImLPR, we conducted a comparison under identical experimental conditions. Since the proposed Patch-InfoNCE loss is incompatible with BEV images, we trained three models: one with the Patch-InfoNCE loss using RIV images, one without this loss using RIV images, and one without this loss using BEV images. We averaged inter-session place recognition results across all sequences in Table. 5, which indicates that RIV representations consistently outperform BEV, even without the Patch-InfoNCE loss. Incorporating the Patch-InfoNCE loss further enhances the patch-level features from RIV images, resulting in improved performance. Visual evidence is presented in Fig. LABEL:fig:overall_dino, which shows the input image alongside an overlay of input reflectivity and DINO features. We apply PCA (PCA) with clipping to distinguish foreground from background. In Fig. LABEL:fig:dino_bev, BEV struggles with feature representation. The yellow box shows distorted DINO features—diagonal in scan 534, linear in scan 1284—while the orange box reveals empty pixels misidentified as foreground, highlighting DINOv2’s limitations in BEV compared to RIV. In Fig. LABEL:fig:dino_riv, RIV shows consistent features between the two images. Notably, white boxes in RIV remain consistent despite a car disappearing, excluding its region. These results indicate that the VFM performs better with RIV representations than with BEV. Additional ablation studies can be found in the appendices.
Table 5: Model Performance with different Input and Loss (AR, AF1: Average Recall and F1 score)
Method Loss Roundabout-O Town-O DCC Roundabout-V Average

### v2

@@ Abstract
LPR (LPR) is a key component in robotic localization, enabling robots to align current scans with prior maps of their environment. While VPR (VPR) has embraced VFM (VFM) to enhance descriptor robustness, LPR has relied on task-specific models with limited use of pre-trained foundation-level knowledge. This is due to the lack of 3D foundation models and the challenges of using VFM (VFM) with LiDAR point clouds. To tackle this, we introduce ImLPR, a novel pipeline that employs a pre-trained DINOv2 VFM to generate rich descriptors for LPR. To the best of our knowledge, ImLPR is the first method to utilize a VFM for LPR while retaining the majority of pre-trained knowledge. ImLPR converts raw point clouds into novel three-channel RIV (RIV) to leverage VFM in the LiDAR domain. It employs MultiConv adapters and Patch-InfoNCE loss for effective feature learning. We validate ImLPR on public datasets and outperform SOTA (SOTA) methods across multiple evaluation metrics in both intra- and inter-session LPR. Comprehensive ablations on key design choices such as channel composition, RIV, adapters, and the patch-level loss quantify each component’s impact. We release ImLPR as open source for the robotics community.
@@ 1 Introduction
Place recognition, encompassing both VPR and LPR, provides an initial localization estimate when seeking to determine if a location has been previously visited, using a prior map of the location. Early learning-based methods relied on specialized architectures trained on domain-specific datasets [1, 2, 3, 4, 5, 6]. While VPR has advanced by using VFMs [7, 8, 9, 10], pre-trained on large-scale datasets [11, 12, 13], LPR has been limited by the lack of suitable 3D foundation models and the modality gap when adapting VFMs to LiDAR data.
Efforts applying foundation models on LiDAR data can be characterized in two ways: developing full 3D foundation models or converting LiDAR data into a format that can be used with VFM. Existing 3D foundation models [14, 15], designed for object detection or indoor scene analysis, are unsuitable for LPR due to their focus on specific tasks and the heavy computation required when working with 3D data with large sets of parameters. On the other hand, converting 3D point clouds into 2D images is a more straightforward approach for leveraging VFM for LPR. Two types of image projection are typically used to create a 2D representation from a 3D point cloud, BEV (BEV) and RIV, as shown in Fig. 1. However, a substantial domain gap exists between natural images used to pre-trained VFMs and projected LiDAR images, particularly when using single-channel [6] or overly large multi-channel inputs [16]. Previous approaches that rely on such representations struggle to effectively leverage pre-trained three-channel vision models.
@@
To tackle these issues, we present ImLPR, a novel pipeline for LPR that adapts DINOv2 [12], a SOTA VFM with robust and transferable feature extraction capabilities. We have designed our system to around a RIV which encodes point clouds using reflectivity, range, and normal ratio. This allows RGB-pretrained VFM to extract discriminative LiDAR features Our empirical results show that each proposed channel has a significant impact on LPR performance, highlighting the importance of channel design. Notably, the three-channel configuration proves more effective for RIV than for BEV. Furthermore, by inserting and training lightweight adapters, we preserve most of the pre-trained DINOv2 weights while adapting it effectively for LPR using RIV input images. Additionally, we propose the Patch-InfoNCE loss, a patch-level contrastive loss, enhancing discriminability and robustness by ensuring consistent feature representations across corresponding RIV patches. Combining LiDAR’s geometric precision and semantic information with DINOv2’s representational power, ImLPR surpasses SOTA LPR methods. Our contributions include:
  • ImLPR is the first LPR pipeline using a VFM while retaining the majority of pre-trained knowledge. Our key innovation lies in a tailored three-channel RIV representation and lightweight convolutional adapters, which seamlessly bridge the 3D LiDAR and 2D vision domain gap. Freezing most DINOv2 layers preserves pre-trained knowledge during training, ensuring strong generalization and outperforming task-specific LPR networks.
@@ 2.2 Vision and 3D Foundation Models for 3D Data
VFMs [12, 13, 11] excel in VPR due to their robust feature representations [7, 9]. In particular, DINOv2 was pre-trained on massive datasets containing 140 million images. In contrast, 3D foundation models like PTv3 [14] and Sonata [15] are trained on smaller datasets with 23-139 thousand point clouds, and focus on object-level tasks, limiting their generality for outdoor LPR. Recently, researchers have attempted to apply VFM to point clouds [25, 26]. However, they focus on object detection or registration, and still rely on visual images, diverging from LPR. Instead, to adapt LiDAR data for use with image-based networks, traditional methods project point clouds into 2D formats like BEV [22] or RIV [16], and then use these images with ImageNet-pretrained CNNs such as ResNet [27]. However, compared to VFM, these methods lack the expressiveness needed for diverse outdoor scenes and are limited by their small training datasets and simple architectures. Additionally, domain differences between LiDAR images and visual images hinder the direct application of VFMs. Fine-tuning the entire model such as LIP-Loc [28], can lead to catastrophic forgetting, reducing generalization by forgetting pre-trained knowledge. To address these challenges, we utilize DINOv2 with a MultiConv adapter to bridge domain differences while preserving almost the entire pre-trained knowledge within the VFM network. We adopt the RIV projection, which captures denser geometric detail and is more similar to normal visual images compared to the sparser BEV images [22, 6].
@@ 3 Methodology
3.1 Overview of ImLPR
ImLPR first renders the scan as an RGB-style image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$ and then computes a global representation $\mathbf{g}\in\mathbb{R}^{t}$ through the mapping function $\Omega=h\circ f$. The VFM-based encoder $f:\mathbb{R}^{H\times W\times 3}\rightarrow\mathbb{R}^{H^{\prime}W^{\prime}\times C}$ ($H^{\prime}=\lfloor H/14\rfloor,\;W^{\prime}=\lfloor W/14\rfloor$) extracts patch features, which are then pooled into the descriptor $\mathbf{g}$ by the aggregator $h:\mathbb{R}^{H^{\prime}W^{\prime}\times C}\rightarrow\mathbb{R}^{t}$. Metric learning optimizes $\Omega$ such that, if the spatial distance $\mathcal{D}(q,i)<\mathcal{D}(q,j)$, then the descriptor distance $d_{g}(q,i)<d_{g}(q,j)$, enforcing this relationship at both the feature and global descriptor levels.
@@ 3.2 Input Data Processing
where $W$ and $H$ denote the image width and height, $r_{i}$ is the range from the origin to $p_{i}$, $f_{\text{up}}$ is the upper bound of the vertical FOV (FOV), and $f$ denotes the total vertical FOV. To further enrich the image with geometric information, we include the three LiDAR channels—reflectivity, range, and normal ratio—in each pixel. Reflectivity offers semantic distinction between objects with different surface properties, while range encodes geometric features of the scene, such as object distance and depth variation. The normal ratio is derived from the singular value ratio of the covariance matrix for the $k$-nearest neighboring points of $p_{i}$. Singular Value Decomposition (SVD) is applied to this covariance matrix, and the logarithmic ratio of the largest to smallest singular values encodes the normal information. This scalar effectively summarizes local geometric variations—such as surface planarity and edge-like structures—into a single channel, offering a computationally efficient alternative to multi-channel normal vectors [16]. This multi-channel image is used as input for ImLPR, as depicted in Fig. 2.
@@ 3.5 Patch-level Contrastive Learning
To encourage learning both locally (patch feature-level) and globally (descriptor-level), we introduce a patch-level contrastive loss in addition to the global objective. This promotes effective gradient flow throughout the network. For patch-level supervision, we mine positive and negative patch pairs from two RIV range images based on geometric consistency.
Positive Patch Mining: We transform both range images into 3D point clouds using their RIV geometry and align them in a common world frame using ground truth poses refined by ICP. Points from the second scan are reprojected onto the first RIV to identify candidate patch correspondences, as illustrated in Fig. 3. Positive pairs are selected based on pixel overlap and range similarity. A patch pair $(p_{1},p_{2})$ is considered positive if the ratio of valid overlapping pixels $\mathcal{P}_{\text{ov}}$, containing range values in both patches, exceeds $\rho_{\text{valid}}=0.5$. For all pixels in $\mathcal{P}_{\text{ov}}$, we compute the per-pixel range differences $d_{i}=r_{1,i}-r_{2,i}$. An adaptive threshold $\tau_{r}$ is computed using the median absolute deviation (MAD): $\tau_{r}=\text{median}(d_{i})+k\cdot\text{MAD}$, ensuring robustness to measurement noise. A pair is confirmed positive if the mean range difference $\Delta_{r}=\frac{1}{|\mathcal{P}_{\text{ov}}|}\sum_{i\in\mathcal{P}_{\text{ov}}}|r_{1,i}-r_{2,i}|$ is less than $\tau_{r}$. This ensures that spatially consistent, range-coherent positive patch pairs are selected.
@@
Negative Patch Mining: To select negative patch pairs, we enforce large vertical and horizontal distances ($v_{\text{dist}}$ and $h_{\text{dist}}$) from the positive patches to avoid spatial overlap. The cylindrical geometry is considered, ensuring that patches near the $0^{\circ}$ and $360^{\circ}$ boundaries are treated as spatially distant.
To supervise patch-level contrastive learning, we propose Patch-InfoNCE, an extension of the InfoNCE loss [13] for local features. Unlike global descriptor losses, Patch-InfoNCE operates on local embeddings from features $\mathbf{F_{1}},\mathbf{F_{2}}\in\mathbb{R}^{H^{\prime}W^{\prime}\times C}$ of two adjacent RIV images. It supports multiple positive patch pairs $(p_{1},p_{2})$ per image and computes their cosine similarity. For each positive pair $F_{1}^{p_{1}}$ and $F_{2}^{p_{2}}$, we select hard negatives: $n_{1}$ from $\mathbf{F_{1}}$ relative to $p_{2}$, and $n_{2}$ from $\mathbf{F_{2}}$ relative to $p_{1}$. The Patch-InfoNCE loss is defined as:
@@
where $\mathcal{P}$ is the set of positive pairs, $\tau_{l}$ is the temperature, and $s_{\text{p}}$, $s_{\text{n},j}$ denote the cosine similarities of positive and negative pairs, respectively. This loss encourages consistent local features at corresponding spatial locations across RIV images, enabling discriminative patch-level embeddings for LPR. The final objective combines both local and global terms as $\mathcal{L}_{\text{final}}=\mathcal{L}_{P}+\lambda\mathcal{L}_{\text{TSAP}}$, where $\lambda$ balances their ratio. Details of $\mathcal{L}_{\text{TSAP}}$ [4] are provided in the appendix. This dual-loss formulation captures both fine-grained context and global semantics, improving overall LPR performance.
@@ 4.1 Experimental Setup
Method Round01-O Round02-O Round03-O Town01-O Town02-O Town03-O Average
@@
(Round: Roundabout, Bold: Best and Underline: Second-Best
Figure 4: Trajectory from two sequences, with red lines marking false positives (fewer red lines indicate better performance). It highlights ImLPR having fewer errors compared to SOTA methods.
@@ 4.2 Intra-session Place Recognition
Scans within 60 seconds of the query are excluded to prevent intra-session matching (i.e., matching within the same trajectory), and processing begins only after 90 seconds to ensure a sufficient database. As shown in Table. 1 and Fig. 4, ImLPR surpasses all baselines on Roundabout-O and Town-O. LoGG3D-Net exhibits lower recall due to the domain shift between training and testing. BEVPlace++ ranks second, capturing scan contours but underperforms ImLPR in narrow alleys and small-scale environments, as ImLPR leverages RIV representations for fine-grained feature extraction. 3D sparse convolution methods like MinkLoc3Dv2 and CASSPR rank third, limited by the absence of large-scale pre-trained models. ImLPR achieves the highest R@1 and F1 score, demonstrating robust and discriminative intra-session LPR performance.
@@ 4.3 Inter-session Place Recognition
Table. 2 and Table. 3 show inter-session results (i.e., between scans collected at different times) for Roundabout-O and Town-O. Each result is shown in the format Database-Query. ImLPR outperforms other methods across both environments, adapting to session-induced variations. BEVPlace++ achieves the second-highest Recall@1, but its lower F1 score indicates reduced precision. We believe this is due to the simple network consisting of 2D convolution without additional mechanisms such as attention. Furthermore, its BEV representation, projecting all data into a single plane, limits its ability to distinguish subtle session-specific variations. As shown in Fig. 5, BEVPlace++ has a lower F1-Recall curve than MinkLoc3Dv2 and CASSPR, despite higher recall. Since the F1-Recall curve captures the trade-off between precision and recall across thresholds, a lower curve suggests a higher false positive rate. MinkLoc3Dv2 and CASSPR follow ImLPR in F1 score with inconsistent results across sequences. This variability highlights the benefit of using large-scale pretrained models for consistent results. LoGG3D-Net ranks lowest due to domain mismatches. ImLPR’s RIV representations precisely capture fine-grained details as shown in Fig. LABEL:fig:dino_riv, and its network design ensures top inter-session LPR performance.
@@
Method Round01-02 Round01-03 Round02-01 Round02-03 Round03-01 Round03-02 Average
@@
Method Town01-02 Town01-03 Town02-01 Town02-03 Town03-01 Town03-02 Average
@@
Table 4: Generalization Assessment
Method DCC NCLT Roundabout-V Average
AR@1 AF1 AR@1 AF1 AR@1 AF1 AR@1 AF1
LoGG3D-Net 0.094 0.214 0.149 0.434 0.131 0.394 0.125 0.347
MinkLoc3Dv2 0.722 0.870 0.622 0.808 0.513 0.764 0.619 0.814
CASSPR 0.683 0.851 0.652 0.810 0.630 0.793 0.655 0.818
BEVPlace++ 0.678 0.839 0.850 0.922 0.655 0.790 0.728 0.850
ImLPR 0.865 0.943 0.834 0.927 0.712 0.852 0.804 0.907
AR@1: Average Recall@1, AF1: Average F1 score
Table 5: Performance Impact of RIV Image Channels
Method Image Channel Roundabout-O Town-O DCC Average
CH1 CH2 CH3 AR@1 AF1 AR@1 AF1 AR@1 AF1 AR@1 AF1
ExpA-1 0.946 0.973 0.956 0.979 0.773 0.884 0.892 0.945
ExpA-2 0.975 0.989 0.980 0.990 0.901 0.951 0.952 0.977
ExpA-3 0.979 0.990 0.985 0.993 0.945 0.970 0.970 0.984
CH1: Reflectivity, CH2: Range, CH3: Normal Ratio
4.4 Ablation Studies
Model Generalization Assessment: To assess the generality of ImLPR, we conducted intra-session LPR on MulRan DCC, NCLT (2012-01-08, 01-15, 01-22), and HeLiPR Roundabout-V without additional training. As these datasets lack a reflectivity channel, normalized intensity was used. For NCLT and Roundabout-V, all methods were validated using accumulated scans as both query and database. The average performance across three sequences is reported in Table. 4, with dataset-specific details provided in the appendix. In DCC, while all methods experienced performance declines, ImLPR demonstrated the most robust performance, excelling in handling sensor differences and temporal shifts. In NCLT and Roundabout-V, it effectively generalized to unseen sensors and environments. Although discrete and noisy intensity from NCLT degraded the semantic quality of RIV images (Fig. LABEL:fig:intensity_dist), ImLPR still attained the highest F1 score by leveraging geometric cues. Unlike competing methods that showed inconsistent rankings across datasets—reflecting sensitivity to specific domains—ImLPR consistently maintained best performance as shown in Fig. 1 and Table. 4. This stability highlights its strong generalization, enabled by RIV and VFM with adapter.
Image Channels: ImLPR uses three image channels—reflectivity, range, and normal ratio—to represent LiDAR point clouds. To assess their individual contributions, we evaluate three configurations: reflectivity only (ExpA-1), reflectivity and range (ExpA-2), and all three channels (ExpA-3), excluding Patch-InfoNCE loss since it relies on the range channel. We average inter-session place recognition results across all sequences, as shown in Table. 5. Each added channel improves AR@1 and AF1. Reflectivity provides a strong semantic baseline, while range adds geometric depth and significantly boosts performance. The normal ratio further improves accuracy by capturing local 3D structure, though its effect is smaller due to the similarity with range, as both channels encode geometric information. These results demonstrate that each channel contributes distinct, complementary features that collectively strengthen the performance of ImLPR.
BEV and RIV Representations with Patch-InfoNCE Loss: We evaluate the effectiveness of BEV and RIV representations using three model variants: BEV without Patch-InfoNCE loss (ExpB-1), RIV without the loss (ExpB-2), and RIV with the loss (ExpB-4), since Patch-InfoNCE is not compatible with BEV. As shown in Table. 6, RIV consistently outperforms BEV, even without Patch-InfoNCE, and further improves with it by enhancing patch-level learning. Visualizations in Fig. LABEL:fig:overall_dino illustrate these differences. BEV struggles with feature representation: the yellow box shows distorted DINO features—diagonal in scan 534 and linear in scan 1284. The orange box highlights empty pixels misidentified as foreground, revealing DINOv2’s limitations. In contrast, RIV produces consistent features across both scans. In Fig. LABEL:fig:dino_riv, white boxes remain stable even as a car disappears in scan 1284, with the region correctly excluded. These results indicate that VFM performs more reliably with RIV, which better retains spatial and semantic patterns.
MultiConv Adapter: Table. 6 shows the benefit of incorporating a MultiConv adapter (ExpB-3) versus not using one (ExpB-4). Even when the same number of last transformer blocks are fine-tuned, the presence of the adapter leads to significant performance improvements, notably increasing both average AR@1 and AF1. This confirms the adapter’s effectiveness in addressing sensor domain shifts while retaining the generalizable representations learned during pre-training. Additional and detailed ablation studies are provided in the appendix.
Table 6: Ablation Study on Input Representation, Loss Function, and Adapter for Inter-Session LPR
Method Input Adapter Loss Roundabout-O Town-O DCC Roundabout-V Average